INAF LBC queue observing, 2013 January 15-16

Observer: M.Faccini, R. Speziali, F.Cusano
Telescope Operator: G.Bechetti
Support Astronomer: H. Khun
Instrument: LBC

Summary

At sunset temperature is about -10, some cirrus far away, wind gust at ~ 15 to 23 m/s
Last nights problems with MODS still not solved
After opening the dome (03:15 UT) we have problems with focus, there is an intra focus on red channel. impossible to solve sending manual offset. Crew have to move the red camera manually.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Giallondo 27D CLJ0956+41 g-sloan, r-sloan 1.5h Total integration time 3h,but 1.5h in the spec for the seeing
Sky flat   g-sloan, r-sloan    
Standard SA98 Us, Y, g, r    

Details

01:00 Turn on LBC

01:05 Error, track error in B_CH, error on array R_CC

01:10 Launch a warm up on LBC

01:35 Turn on LBC, still problems, system rebooted

02:00 LBC now working, we took some bias, the dome is still closed for wind gust

03:15 Dome open

03:19 Start program Battaglia A3-A4

03:20 Slew to the target, LGS3_pointingcheck.ob launched

03:24 dofpia /first X2

03:35 Problem with the focus on the red channel. Intra focus Pupils about 200pixels after manual offset (reaching the limit)

03:47 Dome closed, Technicians are going in the dome to move the red camera. John Hill at the phone suggest an error in the changing position made yesterday.

04:11 Opening dome

04:56 Slewing to a star to check the focusing at the same airmass of the previous pointing

05:11 dopfia first X2

05:19 R_Chan converge after 5 iterations

05:20 Slewing to a star near zenith to double check the improvement

05:22 Double check failed

05:26 Changing offset .....

05:33 We have to move again the red camera (2.5mm)

05:36 Closing dome

06:13 The shutters of the dome do not close due to wind we have to close manually

06:48 Re-opening the dome

06:53 Pointing a star near zenith to check

07:03 dopfia /first /X2

07:14 Seems good, we stop dopfia and slew to the standard field SA98

07:17 dopfia X2

07:23 launched SA98_Usgr.ob, standard photometric field in the filter Us, Y, g and r

07:30 slewing to Giallongo 27D

07:33 dopfia /X2

07:38 pointing ob CLJ0956+41_GR_POINTING.ob for lbcrangebal

07:42 lbcrangebal

07:42 launched CLJ0956+41_GR_a.ob

08:23 dopfia X2

08:31 dopfia failed

08:32 dopfia working again

08:35 dopfia on the pointing ob CLJ0956+41_GR_POINTING.ob

08:44 dopfia launched again

08:54 dopfia

09:00 launched CLJ0956+41_GR_b.ob

09:31 CLJ0956+41_GR_b.ob stopped, images out of focus

09:31 dopfia X2

09:46 dopfia stopped

09:47 dopfia, still problem, it seams that the dopfia is not sending the correction to the z in the red channel

10:05 z changed manually, dopfia sended again

10:16 problem solved, launched CLJ0956+41_GR_b.ob

10:57 launched pointing CLJ0956+41_GR_POINTING.ob

10:58 launched dopfia

11:01 dopfia crashed

11:04 launched CLJ0956+41_GR_c.ob

11:45 dopfia

11:49 launched CLJ0956+41_GR_d.ob

12:31 dopfia

12:36 launched CLJ0956+41_GR_e.ob

13:08 stopped CLJ0956+41_GR_e.ob

13:11 starting test flat field g-sloan r-sloan

13:41 flat filed started g&r-sloan ~22000 counts

13:53 dome closed -- MarcoFaccini - 16 Jan 2013
Topic revision: r4 - 17 Jan 2013, MarcoFaccini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback